# A note on the cohomology of moduli spaces of local shtukas

### v1

@@ 3. Cohomology of moduli spaces of local shtukas
We now apply the constructions of the previous section to the cohomology of mouli space of local shtukas. Let $b\in B(G)$ and let $I$ be a finite set and let $\mu_{\bullet}=(\mu_{i})_{i\in I}$ be a collection of conjugacy classes of cocharacters of $G$, with reflex fields $E_{\bullet}=(E_{i})_{i\in I}$. From the introduction we have the tower $(\operatorname{Sht}_{(G,b,\mu_{\bullet}),K})_{K}$ of moduli spaces of local shtukas. The “intersection cohomology complex” of $\operatorname{Sht}_{(G,b,\mu_{\bullet}),K}$ is the object $f_{K\natural}\mathcal{S}^{\prime}_{W}$ defined before [FS21, Prop. IX.3.2], where $W\in\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})$ is the representation with highest weight $\mu_{\bullet}=(\mu_{i})_{i\in I}$. When $\mu_{\bullet}=\mu$ is a single minuscule cocharacter this is simply the (shifted) compactly supported $\ell$-adic cohomology of the smooth rigid analytic variety $\operatorname{Sht}_{(G,b,\mu),K}$. Let $\rho$ be an irreducible admissible representation of $G_{b}(E)$, with Fargues–Scholze parameter $\varphi$. Consider the $\rho$-isotypic part
@@
where $V\in\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})$ is the dual of the representation with highest weight $\mu_{\bullet}=(\mu_{i})_{i\in I}$, and this is a bounded complex of finite length $G(E)$-representations.
@@
so to understand this better we need to understand the derived stack $\mathfrak{X}_{\varphi}$. A priori, this can be a non-classical stack, and we will have little to say about this case in what follows — to see the correct structure on $R\Gamma(G,b,\mu_{\bullet})[\rho]$ in that case one would need a refinement of Corollary 2.8. We will focus on a particular class of parameters $\varphi$ that turn out to satisfy $\mathfrak{X}_{\varphi}\cong[\ast/S_{\varphi}]$, where $S_{\varphi}$ is the centralizer of $\varphi$, viewed as a point of $\mathfrak{X}$. In what follows, we write $\pi:X^{\Box}\to X$ for the map from $X^{\Box}$ to its GIT quotient $X$. Recall that semisimple $L$-parameters over $\overline{\mathbb{Q}}_{\ell}$ precisely correspond to closed points of $X$.
@@
Let $\varphi$ be a semisimple $L$-parameter over $\overline{\mathbb{Q}}_{\ell}$. Let $\pi^{-1}(\varphi)$ denote the classical (non-derived) scheme-theoretic fibre of $\pi$ at $\varphi$, and let $\pi^{-1}(\varphi)_{red}$ denote its nilreduction.
@@
Concretely, $\varphi$ is strongly semisimple if there are no other $L$-parameters with the same semisimplification as $\varphi$. We note that strongly semisimple $\varphi$ define smooth points of $X^{\Box}$ (or, equivalently, of $\mathfrak{X}$) [^3] This follows from the description of the cotangent complex on $\mathfrak{X}$ [FS21, §VIII.2], using the duality in [FS21, Prop. VIII.2.2].. Now assume that $\varphi$ is generous. Then the underlying classical stack $\mathfrak{X}_{\varphi}^{cl}$ of $\mathfrak{X}_{\varphi}$ is $\pi^{-1}(\varphi)/\widehat{G}$, so by definition $\varphi$ is generous if and only if $\mathfrak{X}_{\varphi}^{cl}\cong[\ast/S_{\varphi}]$. In fact, we will now show that $\mathfrak{X}_{\varphi}$ is classical for generous $\varphi$.
@@
We may prove this component by component on $X$, so let $C\subseteq X$ be a connected component and set $C^{\Box}=\pi^{-1}(C)$; this is connected since $\widehat{G}$ is. By [DHKM24, Thm. 1.8], $C$ is irreducible and reduced. If there are no strongly semisimple parameters on $C$ then the assertion is trivial, so assume that $C_{ss}:=C\cap X_{ss}\neq\emptyset$. Let $C^{\Box}=Y_{1}\cup Y_{2}\cup\dots\cup Y_{r}$ be the decomposition of $C^{\Box}$ into irreducible components. Each component is $\widehat{G}$-invariant, so the sets $\pi(Y_{i})$ are closed. Since $\pi(C^{\Box})=C$, we may without loss of generality assume that $\pi(Y_{1})=C$. Then we must have $C^{\Box}_{ss}:=X^{\Box}_{ss}\cap C^{\Box}\subseteq Y_{1}\setminus(Y_{2}\cup\dots\cup Y_{r})$, since the points in $C^{\Box}_{ss}$ are smooth.
Now set $Y=C\setminus(\pi(Y_{2})\cup\dots\cup\pi(Y_{r}))$ and $Y^{\Box}=\pi^{-1}(Y)$. These are both open and irreducible and we note that $C_{ss}^{\Box}\subseteq Y^{\Box}$ (and $C_{ss}\subseteq Y$). In particular, $Y^{\Box}$ is the GIT quotient of $Y$. Consider the set $U\subseteq Y^{\Box}$ of points whose orbits have maximal dimension; $U$ is open by [New78, Lem. 3.7(c)]. Call this maximal dimension $d^{\prime}$. We let $U^{\prime}\subseteq U$ be subset of points in $U$ whose orbits are closed; this satisfies $U^{\prime}=\pi^{-1}(W^{\prime})$ for some open $W^{\prime}\subseteq Y$ by [New78, Prop. 3.8], and $U^{\prime}\to W^{\prime}$ is a universal geometric quotient. Now consider the set $V\subseteq Y^{\Box}$ of $y\in Y^{\Box}$ for which the dimension of the local ring $\mathcal{O}_{\pi^{-1}(\pi(y)),y}$ of the fibre $\pi^{-1}(\pi(y))$ is minimal; this is an open set by [Mum99, §I.8, Thm. 3 Cor. 3]. Call this minimal dimension $d$. Let $x\in C_{ss}^{\Box}$. By definition of $d^{\prime}$, we must have $d^{\prime}\geq\dim\widehat{G}x$. On the other hand, $\dim\widehat{G}x=\dim\pi^{-1}(\pi(x))$ and $\pi^{-1}(\pi(x))$ is equidimensional by genericity, so we must have $\dim\widehat{G}x\geq d$ by the definition of $d$. In particular, we have $d\leq d^{\prime}$. On the other hand, we have $U\cap V\neq\emptyset$ since $U$ and $V$ are open and $Y^{\Box}$ is irreducible, so choose a point $y\in U\cap V$. Then $d=\dim\mathcal{O}_{\pi^{-1}(\pi(y)),y}$ and $d^{\prime}=\dim\widehat{G}y$ by construction. Since $y\in\widehat{G}y\subseteq\pi^{-1}(\pi(y))$, we must have $d^{\prime}\leq d$, and hence we conclude that $d=d^{\prime}$ and that $C_{ss}\subseteq U$. Since the orbits of semisimple $L$-parameters are closed, we conclude that $C_{ss}^{\Box}\subseteq U^{\prime}$. The converse also holds: Any orbit in $U^{\prime}$ is closed and cannot be in the closure of any other orbit (since such an orbit would need to have dimension bigger than $d^{\prime}$ and lie in $Y^{\Box}$, which is impossible), so we conclude that $C_{ss}^{\Box}=U^{\prime}$ and hence that $W^{\prime}=C_{ss}$. In particular $C_{ss}$ and $C_{ss}^{\Box}$ are open, and $C_{ss}^{\Box}\to C_{ss}$ is a universal geometric quotient. This finishes the proof of the first part of the proposition.
@@
For the second part, note that we have shown that $C_{ss}^{\Box}\to C_{ss}$ is equidimensional (with fibres of dimension $d$) and universally open (since it is a universal geometric quotient), and we know that $C_{ss}$ is reduced since $C$ is. So assume that $\varphi$ is a generous parameter in $C$. Then $\pi^{-1}(\varphi)$ is reduced by assumption, so by [Gro66, Cor. (15.2.3)], it follows that $C_{ss}^{\Box}\to C_{ss}$ is flat in a neighbourhood of $\pi^{-1}(\varphi)$. We may then spread out this neighbourhood using the $\widehat{G}$-action (since we have a geometric quotient) and take the image under $\pi$ to obtain the desired $W$. ∎
@@
If $\varphi$ is a generous $L$-parameter, then $\mathfrak{X}_{\varphi}=\mathfrak{X}_{\varphi}^{cl}=[\ast/\widehat{G}]$. Moreover, for any strongly semisimple $L$-parameter $\varphi$, $\pi$ is flat in a neighbourhood of $\varphi\in X$ if and only if $\varphi\in X$ is smooth.
@@
All strongly semisimple $L$-parameters for $\operatorname{GL}_{n}$ are generous. On the other hand, let $\zeta$ be a primitive $n$-th root of unity and consider the unramified $L$-parameter $\varphi_{n}$ sending Frobenius to $(1,\zeta,\dots,\zeta^{n-1})\in\operatorname{PGL}_{n}(\overline{\mathbb{Q}}_{\ell})$. Then $\varphi_{n}$ is strongly semisimple but not generous, and it is a smooth point in $X$ if and only if $n=2$. We also remark that there is a Zariski open and dense subset of generous parameters inside $X_{ss}$ by [Sta24, Tag 0578] (the full set of generous parameters is locally constructible by [Sta24, Tag 0579], but we do not know if it is open).
@@
By the compatibility between $b$ and $\mu_{\bullet}$, $Z(\widehat{G})^{Q}\subseteq S_{\varphi}$ acts on $V$ via a character $\chi_{b}$. Write $\operatorname{Irr}(S_{\varphi},\chi_{b})$ for the set of irreducible representations of $S_{\varphi}$ on which $Z(\widehat{G})^{Q}$ acts by $\chi_{b}$; this is a finite set. Also write $V_{\varphi}$ for the underlying $W_{E_{\bullet}}$-representation on the image of $V$ under $\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})\to\operatorname{Rep}(S_{\varphi})^{BW_{E_{\bullet}}}$. We can now put everything together to derive the desired formula.
@@
Assume that the Fargues–Scholze parameter $\varphi$ of $\rho$ is generous, and decompose the image of $V$ in $\operatorname{Rep}(S_{\varphi})^{BW_{E_{\bullet}}}$ as $\bigoplus_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}\delta\boxtimes V_{\varphi,\delta}$. Then we have an isomorphism
@@
    We expect that the theorem holds when $\varphi$ is strongly semisimple. For this, one would want to show that $i^{b}_{\ast}\rho$ has a $\operatorname{QCoh}([\ast/S_{\varphi}])$-structure, but this is stronger than $\rho$ having Fargues–Scholze parameter $\varphi$ when $\varphi$ is not generous. In situations when $\varphi$ lifts to a generous parameter on an isogenous group (as in e.g. Remark 3.4), we expect that one can prove this stronger statement, but we have not checked the details.
@@
    The categorical conjecture predicts that $\mathcal{D}^{\varphi}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ is equivalent to $\mathcal{D}(\operatorname{Rep}(S_{\varphi}))$ when $\varphi$ is generous. In particular, each $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ should be a split complex.
@@
    With appropriate modifications, our basic definitions go through with integral coefficients, but making the arguments go through would require more work.
  5. (5)
    Let us briefly compare our results to those of Koshikawa [Kos21], in particular Thm. 1.3 of loc. cit, which says that the $W_{E_{\bullet}}$-representations appearing in $R\Gamma(G,b,\mu_{\bullet})[\rho]$ are subquotients of $V_{\varphi}$, without any assumption on $\varphi$ and allowing integral coefficents. In characteristic 00, we can recover this statement from the assertion that $i^{1\ast}T_{V}i_{b}^{\ast}\rho\in\mathcal{D}^{\varphi}(G(E))^{BW_{E\bullet}}$ (which does not require any assumption on $\varphi$). Moreover, a formula of the form $R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\bigoplus_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}C_{\delta}\boxtimes V_{\varphi,\delta}$ can be deduced straight from the spectral action as in the proof of [Kos21, Thm. 1.3], without the machinery developed here. Indeed, the spectral action shows that the $W_{E_{\bullet}}$-action on $R\Gamma(G,b,\mu_{\bullet})[\rho]$ factors as
    ¯$\overline{\mathbb{Q}}_{\ell}[W_{E_{\bullet}}]\to\operatorname{End}(\mathcal{V})/\mathfrak{m}_{\varphi}\to\operatorname{End}(R\Gamma(G,b,\mu_{\bullet})[\rho]),$
@@
Let $C\subseteq X$ be the connected component containing $\varphi$, let $C^{\Box}$ be its preimage on $X^{\Box}$, and let $Y_{ur}$ the group variety of unramified characters $W_{E}\to Z(\widehat{G})\rtimes Q$. By the local Langlands correspondence for tori, $Y_{ur}$ is isomorphic to the group variety of unramified smooth characters of $G_{b}(E)$. Moreover, $C$ consists of all twists of $\varphi$ by elements of $Y_{ur}$, and all these are generic. In particular, $C^{\Box}\to C$ is smooth with fibres isomorphic to $\widehat{G}/S_{\varphi}$, showing that $\varphi$ is generous.

### v2

@@ 3. Cohomology of moduli spaces of local shtukas
We now apply the constructions of the previous section to the cohomology of moduli space of local shtukas. Let $b\in B(G)$, let $I$ be a finite set and let $\mu_{\bullet}=(\mu_{i})_{i\in I}$ be a collection of conjugacy classes of cocharacters of $G$, with reflex fields $E_{\bullet}=(E_{i})_{i\in I}$. From the introduction we have the tower $(\operatorname{Sht}_{(G,b,\mu_{\bullet}),K})_{K}$ of moduli spaces of local shtukas. The “intersection cohomology complex” of $\operatorname{Sht}_{(G,b,\mu_{\bullet}),K}$ is the object $f_{K\natural}\mathcal{S}^{\prime}_{W}$ defined before [FS21, Prop. IX.3.2], where $W\in\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})$ is the representation with highest weight $\mu_{\bullet}=(\mu_{i})_{i\in I}$. When $\mu_{\bullet}=\mu$ is a single minuscule cocharacter this is simply the (shifted) compactly supported $\ell$-adic cohomology of the smooth rigid analytic variety $\operatorname{Sht}_{(G,b,\mu),K}$. Let $\rho$ be an irreducible admissible representation of $G_{b}(E)$, with Fargues–Scholze parameter $\varphi$. Consider the $\rho$-isotypic part
@@
where $V=W^{\vee}\in\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})$ is the dual of the representation $W$ with highest weight $\mu_{\bullet}=(\mu_{i})_{i\in I}$, and this is a bounded complex of finite length $G(E)$-representations. Note that the restriction of $W$ to the diagonally embedded copy of $Z(\hat{G})^{Q}$ in $\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})$ is necessarily isotypic for some character $\chi$, and the functor $R\Gamma(G,b,\mu_{\bullet})[-]$ will therefore be identically zero unless $\chi$ equals the Kottwitz point of $b$ under the bijection $\pi_{1}(G)_{Q}=Z(\hat{G})^{Q}$, by [Zou22, Lemma 5.3.2]. We will therefore assume this compatibility in the rest of the paper.
@@
so to understand this better we need to understand the derived stack $\mathfrak{X}_{\varphi}$. A priori, this can be a non-classical stack, and we will have little to say about this case in what follows — to see the correct structure on $R\Gamma(G,b,\mu_{\bullet})[\rho]$ in that case one would want a refinement of Corollary 2.9. We will focus on a particular class of parameters $\varphi$ that turn out to satisfy $\mathfrak{X}_{\varphi}\cong[\ast/S_{\varphi}]$, where $S_{\varphi}$ is the centralizer of $\varphi$, viewed as a point of $\mathfrak{X}$. Recall that semisimple $L$-parameters over $\overline{\mathbb{Q}}_{\ell}$ (up to conjugacy) precisely correspond to closed points of $X$. Before defining the classes of parameters that we will be interested in, we state some algebro-geometric conventions that we will use.
Convention 3.1.
When $Y$ is a (classical) scheme of finite type over $\overline{\mathbb{Q}}_{\ell}$, or a disjoint union of such schemes, we will often think of it as its set of $\overline{\mathbb{Q}}_{\ell}$-points (with the Zariski topology) as in classical algebraic geometry. In particular, “the locus of points of $Y$ satisfying a particular property” refers to the subset of $\overline{\mathbb{Q}}_{\ell}$-points with that property. Moreover, $Y_{red}$ will denote the nilreduction of $Y$. Whenever $f:Y_{1}\to Y_{2}$ is a finite type morphism of such schemes and $S\subseteq Y_{2}$ is a subscheme, we write $f^{-1}(S)$ for the (non-derived) fibre product $S\times_{Y_{2}}Y_{1}$. If $S=\{y\}$ is a closed point, we write $f^{-1}(y)$ instead of $f^{-1}(\{y\})$.
We now return to the context of moduli spaces and stacks of $L$-parameters. In what follows, we write $\pi:X^{\Box}\to X$ for the map from $X^{\Box}$ to its GIT quotient $X$.
@@
Let $\varphi$ be a semisimple $L$-parameter over $\overline{\mathbb{Q}}_{\ell}$.
@@
Concretely, $\varphi$ is strongly semisimple if there are no other $L$-parameters with the same semisimplification as $\varphi$. We have the following lemma:
Lemma 3.3.
Let $\varphi$ be a strongly semisimple $L$-parameter. Then $\varphi$ defines a smooth point of $X^{\Box}$ or, equivalently, of $\mathfrak{X}$.
Proof.
By [FS21, §VIII.2], the tangent complex of $\mathfrak{X}$ at $\varphi$ is $R\Gamma(W_{E},\mathrm{ad}(\varphi))[1]$, which is concentrated in [1,1][-1,1], and the obstruction group is $H^{2}(W_{E},\mathrm{ad}(\varphi))$. Thus, it suffices to prove that $H^{2}(W_{E},\mathrm{ad}(\varphi))=0$. By [FS21, Prop. VIII.2.2], we have
$H^{2}(W_{E},\mathrm{ad}(\varphi))\cong H^{0}(W_{E},\mathrm{ad}(\varphi)(1))^{\ast},$
[…]
For any semisimple $L$-parameter $\varphi$, the underlying classical stack $\mathfrak{X}_{\varphi}^{cl}$ of $\mathfrak{X}_{\varphi}$ is $\pi^{-1}(\varphi)/\widehat{G}$], so by definition $\varphi$ is generous if and only if $\mathfrak{X}_{\varphi}^{cl}\cong[\ast/S_{\varphi}]$. In fact, we will now show that $\mathfrak{X}_{\varphi}$ is classical for generous $\varphi$. Recall our conventions from Convention 3.1. We isolate a general geometric invariant theory result from the argument. For this, we will need some basic notions and results from affine geometric invariant theory. We refer the reader to [MFK94, §1.2] and [New78, Thm. 3.5] for the basic properties that we will use.
Lemma 3.4.
Let $S$ be an irreducible affine $\overline{\mathbb{Q}}_{\ell}$-variety with an action of $\widehat{G}$ and let $T=S\sslash\widehat{G}$ be its GIT quotient. Write $f:S\to T$ for the quotient map. Let $U\subseteq S$ denote the union of the $\widehat{G}-$orbits which are closed and not contained in the closure of any other $\widehat{G}-$orbit. Then there exists an open subset $V\subseteq T$ such that $U=f^{-1}(V)$, and $f:U\to V$ is a universal geometric quotient.
Proof.
Consider the set $U_{1}\subseteq S$ of points whose orbits have maximal dimension, and denote this maximal dimension by $d^{\prime}$. By [New78, Lem. 3.7(c)], $U_{1}$ is open. Next, let $U_{2}\subseteq U_{1}$ be the subset of points whose orbits are closed. By [New78, Prop. 3.8], $U_{2}=f^{-1}(V)$ for some open $V\subseteq T$ , and $U_{2}\to V$ is a universal geometric quotient [^3] For the reader looking at this reference, we note that the notions of orbit spaces and geometric quotients, and their relation, is covered in [New78, p. 31, Def.] and [New78, p. 57, Def.], and that geometric quotients in characteristic 0 in affine GIT are universal [MFK94, Amplification 1.3].. Thus, it suffices to show that $U=U_{2}$. To show that $U_{2}\subseteq U$, note that any orbit in $U_{2}$ is closed and cannot be in the closure of any other orbit, since such an orbit would need to have bigger dimension, which is impossible by the definition of $U_{2}$.
It therefore remains to show that $U\subseteq U_{2}$. Consider the set $U_{3}\subseteq S$ of $x\in S$ for which the dimension of the local ring $\mathcal{O}_{f^{-1}(f(x)),x}$ of the fibre $f^{-1}(f(x))$ is minimal. By [Mum99, §I.8, Thm. 3 Cor. 3], $U_{3}$ is open. Call this minimal dimension $d$. Let $x\in U$. By definition of $d^{\prime}$, we must have $d^{\prime}\geq\dim\widehat{G}x$. On the other hand, $\widehat{G}x=f^{-1}(f(x))$ by the definition of $U$ and it is equidimensional (since it is a single orbit), so we must have $\dim\widehat{G}x\geq d$ by the definition of $d$. In particular, we have $d\leq d^{\prime}$. On the other hand, we have $U_{1}\cap U_{3}\neq\emptyset$ since $U_{1}$ and $U_{3}$ are open and non-empty, and $S$ is irreducible. Choose a point $y\in U_{1}\cap U_{3}$. Then $d=\dim\mathcal{O}_{f^{-1}(f(y)),y}$ and $d^{\prime}=\dim\widehat{G}y$ by definition. Since $\widehat{G}y\subseteq f^{-1}(f(y))$ and $\widehat{G}y$ is equidimensional, we must have $d^{\prime}\leq d$, and hence we conclude that $d=d^{\prime}$ and hence that $U\subseteq U_{1}$ (since we proved that $d\leq\dim\widehat{G}x\leq d^{\prime}$ for $x\in U$). Since the orbits in $U$ are closed, we conclude that $U\subseteq U_{2}$, as desired, finishing the proof. ∎
@@
We start with the first part. As a first reduction step, note that we may prove the proposition component by component on $X$, so let $C\subseteq X$ be a connected component and set $C^{\Box}=\pi^{-1}(C)$; this is connected since $\widehat{G}$ is. Moreover, if there are no strongly semisimple parameters on $C$ then the assertion is trivial, so assume that $C_{ss}:=C\cap X_{ss}\neq\emptyset$. By [DHKM24, Thm. 1.7], $C$ is irreducible and reduced. However, $C^{\Box}$ might not be irreducible, preventing us from applying Lemma 3.4 directly (note that the condition on the orbits included in $U$ in the notation of Lemma 3.4 precisely corresponds to the definition of strong semisimplicity). Instead, we will show that the Proposition follows by applying Lemma 3.4 to a particular irreducible component of $C^{\Box}$.
Write $C^{\Box}=Y_{1}\cup Y_{2}\cup\dots\cup Y_{r}$ as the union of its irreducible components. Each component is $\widehat{G}-$invariant, so the sets $\pi(Y_{i})$ are closed. Since $\pi(C^{\Box})=C$, we may without loss of generality assume that $\pi(Y_{1})=C$. By exactness of $\widehat{G}-$invariants, it follows that $C$ is the GIT quotient of $Y_{1}$ as well. If $C_{ss}^{\Box}:=\pi^{-1}(C_{ss})\subseteq Y_{1}$, then we can apply Lemma 3.4 to conclude the first part of the Proposition. To see that $C_{ss}^{\Box}\subseteq Y_{1}$, note that each orbit in $C_{ss}^{\Box}$ has to intersect $Y_{1}$ since the orbits are closed and $C_{ss}\subseteq\pi(Y_{1})$, and then use $\widehat{G}-$invariance of $Y_{1}$ [^4] In fact, since the points in $C_{ss}^{\Box}$ are smooth by Lemma 3.3, they only lie on $Y_{1}$. This means that $Y_{1}$ is the only component of $C^{\Box}$ mapping surjectively onto $C$.. This finishes the proof of the first part.
@@
For the second part, note that we have shown that $C_{ss}^{\Box}\to C_{ss}$ is equidimensional and universally open (since it is a universal geometric quotient), and we know that $C_{ss}$ is reduced since $C$ is. So assume that $\varphi$ is a generous parameter in $C$. Then $\pi^{-1}(\varphi)$ is reduced by assumption, so by [Gro66, Cor. (15.2.3)], it follows that $C_{ss}^{\Box}\to C_{ss}$ is flat in a neighbourhood of $\pi^{-1}(\varphi)$. We may then spread out this neighbourhood using the $\widehat{G}-$action (since we have a geometric quotient) and take the image under $\pi$ to obtain the desired $W$. ∎
@@
If $\varphi$ is a generous $L$-parameter, then $\mathfrak{X}_{\varphi}=\mathfrak{X}_{\varphi}^{cl}=[\ast/S_{\varphi}]$. Moreover, for any strongly semisimple $L$-parameter $\varphi$, $\pi$ is flat in a neighbourhood of $\varphi\in X$ if and only if $\varphi\in X$ is smooth.
@@
Let us make some remarks on the prevalence of generous parameters. Here, we think of $L$-parameters as Weil–Deligne parameters.
  1. (1)
    In general, there is a Zariski open and dense subset of generous parameters inside $X_{ss}$ by [Sta24, Tag 0578] (the full set of generous parameters is locally constructible by [Sta24, Tag 0579], but we do not know if it is open).
  2. (2)
    Let us indicate a case when strongly semisimple and generous parameters can be described rather explicitly. Assume that $G$ is split, and consider the unramified (or principal) component $\mathfrak{X}_{ur}$ of $\mathfrak{X}$, consisiting of parameters $(\varphi,N)$ for which $\varphi$ is unramified (and hence equivalent to a single element in $\widehat{G}$), which is studied in detail in [Hel23, §2]. Write $X^{\Box}_{ur}$ and $X_{ur}$ for the corresponding components of $X^{\Box}$ and $X$, respectively. As noted in a footnote in the proof of Proposition 3.5, $\mathfrak{X}_{ur}$ has a unique irreducible component mapping surjectively onto the GIT quotient, and it contains all strongly semisimple parameters. In this case, it is the component where $N=0$, hence isomorphic to $\widehat{G}/\widehat{G}$] (the action being by conjugation). The strongly semisimple parameters correspond to those regular semisimple elements in $\widehat{G}$ which do not lie on other components of $\mathfrak{X}_{ur}$. By [Ste65, 6.4, Cor.], we then have
    (2) $X_{ur}^{\Box}=\widehat{G}\to X_{ur}=\widehat{G}\sslash\widehat{G}\cong\widehat{T}\sslash N(\widehat{T})=\widehat{T}\sslash W,$
    where $\widehat{T}$ is a maximal torus of $\widehat{G}$, with normalizer N(T^)N($\widehat${T}), and $W=N(\widehat{T})/\widehat{T}$ is the Weyl group. It is well known in the theory of algebraic groups how this map can be analysed, but let us sketch one way. The morphism $\widehat{T}/N(\widehat{T})]\to[\widehat{G}/\widehat{G}$] of quotient stacks (induced by $\widehat{T}\subseteq\widehat{G}$) is an isomorphism over the regular semisimple locus (on both sides). In particular, in the smooth topology and over the regular semisimple locus, the fibres of the map in equation (2) are equivalent to the fibres of $\widehat{T}\to\widehat{T}\sslash W$. The latter are smooth if and only if they are étale, i.e. when the centralizer of the regular semisimple element is $\widehat{T}$ itself. In particular, this is always the case when the derived group of $\widehat{G}$ is simply connected [Ste65, 2.10, Rem.]. In that case, all strongly semisimple unramified parameters are generous.
[…]
  3. (3)
    The previous part of the remark shows that all unramified strongly semisimple $L$-parameters for $\operatorname{GL}_{n}$ are generous. In fact, one may bootstrap this to show that all strongly semisimple $L$-parameters for $\operatorname{GL}_{n}$ are generous. This follows from the fact that any component of $\mathfrak{X}$ (for $\operatorname{GL}_{n}$) is isomorphic to a product of unramified components for some $\operatorname{GL}_{d}$’s (possibly for an extension of $F$). For a reference for this, see the very end of [BZCHN24, §5.2].
@@
By the compatibility between $b$ and $\mu_{\bullet}$ (as recalled in the beginning of this section), $Z(\widehat{G})^{Q}\subseteq S_{\varphi}$ acts on $V$ via a character $\chi_{b}$. Write $\operatorname{Irr}(S_{\varphi},\chi_{b})$ for the set of irreducible representations of $S_{\varphi}$ on which $Z(\widehat{G})^{Q}$ acts by $\chi_{b}$. Also write $V_{\varphi}$ for the image of $V$ under $\operatorname{Rep}(\widehat{G}^{I}\rtimes Q_{\bullet})\to\operatorname{Rep}(S_{\varphi})^{BW_{E_{\bullet}}}$. We can now put everything together to derive the desired formula.
@@
Let $\rho$ be an irreducible admissible representation of $G_{b}(E)$ and assume that its Fargues–Scholze parameter $\varphi$ is generous. Decompose $V_{\varphi}$ in $\operatorname{Rep}(S_{\varphi})^{BW_{E_{\bullet}}}$ as $\bigoplus_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}\delta\boxtimes V_{\varphi,\delta}$. Then we have an isomorphism
@@
    We expect that the theorem holds when $\varphi$ is strongly semisimple. For this, one would want to show that $i^{b}_{\ast}\rho$ has a $\operatorname{QCoh}([\ast/S_{\varphi}])$-structure, but this is stronger than $\rho$ having Fargues–Scholze parameter $\varphi$ when $\varphi$ is not generous. In situations when $\varphi$ lifts to a generous parameter on an isogenous group, we expect that one can prove this stronger statement, but we have not checked the details.
@@
    The categorical conjecture [FS21, Conj. X.1.4] predicts that $\mathcal{D}^{\varphi}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ is equivalent to $\mathcal{D}(\operatorname{Rep}(S_{\varphi}))$ when $\varphi$ is generous. In particular, each $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ should be a split complex.
@@
    Let us briefly compare our results to those of Koshikawa [Kos21], in particular Thm. 1.3 of loc. cit, which says that the $W_{E_{\bullet}}$-representations appearing in $R\Gamma(G,b,\mu_{\bullet})[\rho]$ are subquotients of $V_{\varphi}$, without any assumption on $\varphi$ and allowing integral coefficents. In characteristic 0, we can recover this statement from the assertion that $i^{1\ast}T_{V}i_{b}^{\ast}\rho\in\mathcal{D}^{\varphi}(G(E))^{BW_{E\bullet}}$ (which does not require any assumption on $\varphi$). Moreover, a formula of the form $R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\bigoplus_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}C_{\delta}\boxtimes V_{\varphi,\delta}$ can be deduced straight from the spectral action as in the proof of [Kos21, Thm. 1.3], without the machinery developed here. Indeed, the spectral action shows that the $W_{E_{\bullet}}$-action on $R\Gamma(G,b,\mu_{\bullet})[\rho]$ factors as
    $\overline{\mathbb{Q}}_{\ell}[W_{E_{\bullet}}]\to\operatorname{End}(\mathcal{V})/\mathfrak{m}_{\varphi}\to\operatorname{End}(R\Gamma(G,b,\mu_{\bullet})[\rho]),$
@@
Let $C\subseteq X$ be the connected component containing $\varphi$, let $C^{\Box}$ be its preimage in $X^{\Box}$, and let $Y_{ur}$ the group variety of unramified characters $W_{E}\to Z(\widehat{G})\rtimes Q$. By the local Langlands correspondence for tori, $Y_{ur}$ is isomorphic to the group variety of unramified smooth characters of $G_{b}(E)$. As remarked just after [FS21, Def X.1.2] [^6] The deformation-theoretic argument in [FS21] appears to be incomplete. It shows that the family of unramified twists of $\varphi$ is an open subset of $C$, but one needs an argument to show that it is closed. If $Z(\widehat{G})^{Q}$ is finite, i.e. the connected split center of $G$ is trivial, then there are no non-trivial unramified characters, hence the family is a point and therefore closed. In general, one can reduce to this case by considering the map from the moduli space of $L$-parameters of $G$ to the moduli space of $L$-parameters of the quotient of $G$ by its connected split center. The description of $C$ can also be extracted from [DHKM24, Thm. 1.7]., $C$ consists of all twists of $\varphi$ by elements of $Y_{ur}$, and all these are strongly semisimple. In particular, $C^{\Box}\to C$ is smooth with fibres isomorphic to $\widehat{G}/S_{\varphi}$, showing that $\varphi$ is generous.